OpenStudy (anonymous):

(1 1/2) to the 2 power - (4-2 1/3)

OpenStudy (anonymous):

This? \[\left(1\frac{1}{2}\right)^{2}-\left(4-2\frac{1}{3}\right)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\left(\frac{3}{2}\right)^{2} - \left(\frac{12}{3} - \frac{7}{3}\right)\]\[\frac{9}{4} - \frac{5}{3}\]\[\frac{27}{12} - \frac{20}{12} = \frac{7}{12}\]

OpenStudy (anonymous):

thank u